OpenStudy (anonymous):

The sum of 2 consecutive odd integers is 92. Find the numbers. If x represents the smaller integer, then the larger integer is x + 1 x + 2 2x + 1

OpenStudy (anonymous):

X+(X+2)=92 2X+2=92 2X=90 X=45 AND (X+2)=45+2=47 SO 45 AND 47 ARE THE NUMBERS

OpenStudy (anonymous):

IF S IS SMALLER THEN X+2 IS LARGER

OpenStudy (anonymous):

thankx that helped

OpenStudy (anonymous):

WELCOME!!!